OpenStudy (priyar):

Let f(x) be a non-negative continuous function such that the area bounded by the curve y=f(x),x-axis and the ordinates x=pi/4 and..(the rest of the q is below)

OpenStudy (priyar):

and x=\[\beta>\pi/4 \] is (\[(\beta \sin(\beta)+\pi/4 \cos (\beta)+\sqrt{2}\beta )\]).Then f(pi/2) is?

OpenStudy (priyar):

@hartnn

OpenStudy (priyar):

@zepdrix @Astrophysics

OpenStudy (anonymous):

I can't help here but try using desmos calculator! it helped me a lot in the past

ganeshie8 (ganeshie8):

Hint : \[f(\beta) = \dfrac{d}{d\beta}\int\limits_{\pi/4}^{\beta} f(x)\,dx \]

OpenStudy (priyar):

thanks for the hint @ganeshie8 ..I got it!

OpenStudy (priyar):

the answer is 1-pi/4 +(2)^1/2 !

OpenStudy (anonymous):

good questions :)